OpenStudy (anonymous):

please explain how I can solve -k - 16 + 4k -9 and please don't just do the problem for me.If you do I won't understand it and won't be able to do my other problems...

OpenStudy (anonymous):

The whole idea of these problems is to solve for the variable, or isolate k in this.

OpenStudy (anonymous):

What you do to one side you must do to the other, Example) 4k = 2 divide 4 from each side, you do this because you want to cancel the 4 with the k (isolate it), so you'd get k = 2/4, k = 1/2

OpenStudy (anonymous):

Hope that helps.

OpenStudy (anonymous):

You seem to be missing an equals sign.

OpenStudy (doc.brown):

@iambatman and @douglaswinslowcooper have great points. You can't do any balancing without an equal sign. Do you mean -k-16 = 4k-9 ?